# WeTok: Powerful Discrete Tokenization for High-Fidelity Visual Reconstruction

### v1

@@ Abstract
Visual tokenizer is a critical component for vision generation. However, the existing tokenizers often face unsatisfactory trade-off between compression ratios and reconstruction fidelity. To fill this gap, we introduce a powerful and concise WeTok tokenizer, which surpasses the previous leading tokenizers via two core innovations. (1) Group-wise lookup-free Quantization (GQ). We partition the latent features into groups, and perform lookup-free quantization for each group. As a result, GQ can efficiently overcome memory and computation limitations of prior tokenizers, while achieving a reconstruction breakthrough with more scalable codebooks. (2) Generative Decoding (GD). Different from prior tokenizers, we introduce a generative decoder with a prior of extra noise variable. In this case, GD can probabilistically model the distribution of visual data conditioned on discrete tokens, allowing WeTok to reconstruct visual details, especially at high compression ratios. Extensive experiments on mainstream benchmarks show superior performance of our WeTok. On the ImageNet 50k validation set, WeTok achieves a record-low zero-shot rFID (WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ vs. SD-VAE 3.5: $\mathbf{0.19}$). Furthermore, our highest compression model achieves a zero-shot rFID of 3.49 with a compression ratio of 768, outperforming Cosmos (384) $4.57$ which has only 50% compression rate of ours. Code and models are available: https://github.com/zhuangshaobin/WeTok.
@@ 1 Introduction
Finally, we conduct extensive experiments on mainstream benchmarks, via scaling WeTok across group size, model size, and training data size. Moreover, we pre-train our WeTok on a 400M general-domain dataset across multiple compression ratios. As illustrated in Fig. 1 (a), WeTok consistently outperforms the state-of-the-art continuous and discrete tokenizers, e.g., rFID on ImageNet 50k validation set: WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ (Batifol et al., 2025) vs. SD-VAE 3.5: $\mathbf{0.19}$ (Esser et al., 2024a). Furthermore, our highest compression model also achieves the superior reconstruction performance, e.g., rFID on ImageNet 50k validation set: WeTok: $3.59$ vs. Cosmos: $4.57$ (Agarwal et al., 2025), while Cosmos ($384$) only has 50% compression ratio of our WeTok ($768$), showing effectiveness and efficiency of WeTok.
@@ 3.1 Preliminaries
Vector Quantized Variational Autoencoder (Esser et al., 2020). Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQVAE first compresses it into latent feature $\mathcal{U}=\mathcal{E}(\mathcal{I})$, $\mathcal{U}\in\mathbb{R}^{h\times w\times d}$, through the encoder $\mathcal{E}$. Then it is quantized into latent codes $\mathcal{Q}$ by searching the nearest neighbor in the codebook $\mathcal{C}\in\mathbb{R}^{K\times d}$,
$\mathcal{Q}[i,j]=\mathop{\arg\min}_{\mathbf{c}_{k}\in{\mathcal{C}}}\|\mathcal{U}[i,j]-\mathbf{c}_{k}\|^{2},$ (1)
$\mathcal{U}_{\mathcal{Q}}=\mathcal{U}+\text{sg}[\mathcal{Q}-\mathcal{U}],$ (2)
where $\text{sg}[\cdot]$ is stop-gradient operation. Finally, it is reconstructed into image space $\hat{\mathcal{I}}=\mathcal{G}(\mathcal{U}_{\mathcal{Q}})$ through the decoder $\mathcal{G}$. The loss function consists of five parts,
@@
Lookup-Free Quantization (Yu et al., 2024a). LFQ introduces an implicit and non-learnable codebook $\mathcal{C}_{\text{LFQ}}=\{-1,1\}^{d}$ to perform lookup-free quantization on each channel of latent feature,
@@
Since the codebook in LFQ is fixed, there is no need for codebook loss during LFQ training. To address the issue of codebook utilization collapse in VQVAE, LFQ introduces entropy loss,
$\mathcal{L}_{\text{Entropy}}(\mathcal{U})=\underbrace{\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}H(q(\mathbf{c}|\mathcal{U}[i,j]))}_{\text{Token Entropy Loss}}-\zeta\underbrace{H(\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))}_{\text{Codebook Entropy Loss}}.$ (5)
Binary Spherical Quantization (Zhao et al., 2024b). When increasing the codebook size, i.e., increasing $d$, the $H(\cdot)$ calculation in the $\mathcal{L}_{\text{entropy}}$ of LFQ will lead to significant memory consumption. To alleviate this issue, BSQ first rewrites the token entropy loss as $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Next, BSQ rewrites the codebook entropy loss as $\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$ by assuming the approximation $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))\approx\prod_{k=1}^{d}\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Both operations reduce the variable space of the $H(\cdot)$ calculation from $\{-1,1\}^{d}$ to the linear combination of $d$ $\{-1,1\}$, significantly decreasing memory consumption. However, due to the approximation proposed by BSQ for the codebook entropy loss, the entropy loss calculation in BSQ introduces errors. As a result, under the same codebook size, the performance of BSQ is inferior to that of LFQ.
@@ 3.2 Group-Wise Lookup-Free Quantization
$\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j])\approx\prod_{k=1}^{g}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k]).$ (8)
@@ 3.3 Generative Decoder
To ensure training stability, we employ a two-stage training process. In the first stage, we train our WeTok with the reconstruction loss, i.e., Eq. 3, 7 and 9. In the second stage, we adapt the model for generative tasks. Specifically, we expand the channel dimension of the conv_in layer in decoder to accept z as additional input. To preserve the powerful reconstruction capabilities learned in the first stage, the newly added weights for these channels are zero-initialized. This strategy ensures that at the beginning of the second stage, the decoder’s behavior is identical to its pre-trained state. Consequently, this versatile approach allows our GD to be effectively integrated with various pre-trained discrete tokenizers that share a similar architectural foundation.
@@ 4 Experiments
SD-VAE 3.5 (Esser et al., 2024b) Continuous - 8 24 1.66 31.08 0.90 0.19 31.19 0.90
FLUX-VAE (Labs, 2024) Continuous - 8 24 1.35 32.32 0.93 0.18 32.74 0.92

### v2

@@ Abstract
Visual tokenizer is a critical component for vision generation. However, the existing tokenizers often face unsatisfactory trade-off between compression ratios and reconstruction fidelity. To fill this gap, we introduce a powerful and concise WeTok tokenizer, which surpasses the previous leading tokenizers via two core innovations. (1) Group-wise lookup-free Quantization (GQ). We partition the latent features into groups, and perform lookup-free quantization for each group. As a result, GQ can efficiently overcome memory and computation limitations of prior tokenizers, while achieving a reconstruction breakthrough with more scalable codebooks. (2) Generative Decoding (GD). Different from prior tokenizers, we introduce a generative decoder with a prior of extra noise variable. In this case, GD can probabilistically model the distribution of visual data conditioned on discrete tokens, allowing WeTok to reconstruct visual details, especially at high compression ratios. Extensive experiments on mainstream benchmarks show superior performance of our WeTok. On the ImageNet 50k validation set, WeTok achieves a record-low zero-shot rFID (WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ vs. SD-VAE 3.5: $\mathbf{0.19}$) with a $\mathbf{400\%}$ compression ratio. Furthermore, our highest compression model achieves a zero-shot rFID of $3.49$ with a compression ratio of 768, outperforming Cosmos (384) $4.57$ which has only 50% compression rate of ours. Code and models are available: https://github.com/zhuangshaobin/WeTok.
@@ 1 Introduction
Finally, we conduct extensive experiments on mainstream benchmarks, via scaling WeTok across group size, model size, and training data size. Moreover, we pre-train our WeTok on a 400M general-domain dataset across multiple compression ratios. As illustrated in Fig. 1 (a), WeTok consistently outperforms the state-of-the-art continuous and discrete tokenizers with a $\mathbf{400\%}$ compression ratio, e.g., rFID on ImageNet 50k validation set: WeTok: $\mathbf{0.12}$ vs. FLUX-VAE: $\mathbf{0.18}$ (Batifol et al., 2025) vs. SD-VAE 3.5: $\mathbf{0.19}$ (Esser et al., 2024a). Furthermore, our highest compression model also achieves the superior reconstruction performance, e.g., rFID on ImageNet 50k validation set: WeTok: $3.59$ vs. Cosmos: $4.57$ (Agarwal et al., 2025), while Cosmos ($384$) only has 50% compression ratio of our WeTok ($768$), showing effectiveness and efficiency of WeTok.
@@ 3.1 Preliminaries
Vector Quantized Variational Autoencoder (Esser et al., 2020). Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQVAE first compresses it into latent feature $\mathcal{U}=\mathcal{E}(\mathcal{I})$, $\mathcal{U}\in\mathbb{R}^{h\times w\times d}$, through the encoder $\mathcal{E}$. Then it is quantized into latent codes $\mathcal{Q}$ by searching the nearest neighbor in the codebook $\mathcal{C}\in\mathbb{R}^{K\times d}=[\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{K}]^{\top}$,
$\mathcal{Q}[i,j]=\mathop{\arg\min}_{\mathbf{c}\in{\mathcal{C}}}\|\mathcal{U}[i,j]-\mathbf{c}\|^{2},$ (1)
Notice that here we define $\mathcal{Q}$ to have a backward gradient only with respect to $\mathcal{C}$, and introduce $\mathcal{U}_{\mathcal{Q}}=\mathcal{U}+\text{sg}[\mathcal{Q}-\mathcal{U}]$ as a variable that shares the same value as $\mathcal{Q}$ but has a backward gradient only with respect to $\mathcal{U}$. Here, $\text{sg}[\cdot]$ denotes the stop-gradient operation.
Finally, the intermediate quantized result $\mathcal{Q}$ is reconstructed into image space $\hat{\mathcal{I}}=\mathcal{G}(\mathcal{U}_{\mathcal{Q}})$ through the decoder $\mathcal{G}$. The loss function of VQ-VAE consists of the following five parts,
@@
Lookup-Free Quantization (Yu et al., 2024a). LFQ introduces an implicit and learning-free codebook $\mathcal{C}_{\text{LFQ}}=\{-1,1\}^{d}$ to perform lookup-free quantization on each channel of latent feature,
@@
Since the codebook in LFQ is fixed, there is no need for codebook loss during LFQ training. To address the issue of codebook utilization collapse in VQVAE, LFQ introduces the following entropy loss in replace of the commitment loss in equation 2,
$\mathcal{L}_{\text{Entropy}}(\mathcal{U})=\underbrace{\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}H\left(q(\mathbf{c}|\mathcal{U}[i,j])\right)}_{\text{Token Entropy Loss}}-\zeta\underbrace{H\left(\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j])\right)}_{\text{Codebook Entropy Loss}}.$ (4)
where $H(X)=-\sum_{Xi\in X}X_{i}\text{log}X_{i}$ , $q(\mathbf{c}|\mathcal{U}[i,j])$ denote the conditional distribution of $\mathbf{c}$ given $\mathcal{U}[i,j]$ , and $\zeta$ refers to the weight of codebook entropy loss.
Binary Spherical Quantization (Zhao et al., 2024b). When increasing the codebook size, i.e., increasing $d$, the $H(\cdot)$ calculation in the $\mathcal{L}_{\text{Entropy}}$ of LFQ will lead to significant memory consumption. To alleviate this issue, BSQ seeks to handle each binary of LFQ seperately. BSQ first rewrites the token entropy loss into binary seperated form as $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Next, BSQ rewrites the codebook entropy loss as $\sum_{k=1}^{d}H(q_{B}(\mathbf{c}[k]|\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}\mathcal{U}[i,j,k]))$ by assuming the approximation $\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j]))\approx\prod_{k=1}^{d}\frac{1}{hw}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{B}(\mathbf{c}[k]|\mathcal{U}[i,j,k]))$. Both operations reduce the variable space of the $H(\cdot)$ calculation from $\{-1,1\}^{d}$ to the linear combination of $d$ $\{-1,1\}$, significantly decreasing memory consumption. However, as the binary approximation for the codebook entropy loss proposed by BSQ would introduce additional errors, the performance of BSQ is inferior to that of LFQ under the same codebook size.
@@ 3.2 Group-Wise Lookup-Free Quantization
$\sum_{i=1}^{h}\sum_{j=1}^{w}q(\mathbf{c}|\mathcal{U}[i,j])=\sum_{i=1}^{h}\sum_{j=1}^{w}\prod_{k=1}^{g}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k])\approx\prod_{k=1}^{g}\sum_{i=1}^{h}\sum_{j=1}^{w}q_{\text{G}}(\mathbf{c}_{k}|\mathcal{U}_{\text{G}}[i,j,k]).$ (7)
@@ 3.3 Generative Decoder
To ensure training stability, we employ a two-stage training process. In the first stage, we train our WeTok with the reconstruction loss, i.e., Eq. 2, 6 and 8. In the second stage, we adapt the model for generative tasks. Specifically, we expand the channel dimension of the conv_in layer in decoder to accept z as additional input. To preserve the powerful reconstruction capabilities learned in the first stage, the newly added weights for these channels are zero-initialized. This strategy ensures that at the beginning of the second stage, the decoder’s behavior is identical to its pre-trained state. Consequently, this versatile approach allows our GD to be effectively integrated with various pre-trained discrete tokenizers that share a similar architectural foundation.
4 Experiments
@@
SD-VAE 3.5 (Esser et al., 2024b) Continuous - 8 6 1.66 31.08 0.90 0.19 31.19 0.90
FLUX-VAE (Labs, 2024) Continuous - 8 6 1.35 32.32 0.93 0.18 32.74 0.92